OpenStudy (anonymous):

for a function f(x) to have an inverse , than the functions derivative should be greater than 0, f(x)'>0, right? so if f(x)=1/x, whose derivative is f(x)'= -1/x^2, than f(x) doesn't have an inverse

OpenStudy (anonymous):

@Kainui

OpenStudy (perl):

the derivative has to be monotone

ganeshie8 (ganeshie8):

do you remember that vertical stick test ?

ganeshie8 (ganeshie8):

it seems f(x) = 1/x passes that test right ?

OpenStudy (anonymous):

Okay, I try.. :) Thanks.. :)

OpenStudy (jhannybean):

and include where the derivative does not exist?

OpenStudy (perl):

the derivative has to be either f ' (x) >= 0 or f '(x) <= 0 on the entire interval

OpenStudy (kainui):

Honestly I am more a physicist/chemist so I will tell you that it has an inverse and that it is its own inverse since in all practicality it will give you exactly the inverse. Just as a little bit of 'proof' look: \[\LARGE f(f^{-1}(x))=x\] this is the definition of an inverse as far as I'm concerned. \[\LARGE f(x)=\frac{1}{x}\] this is our function. \[\LARGE f(f(x))=\frac{1}{\frac{1}{x}}=x\] Making the distinction here is sort of silly to me I guess. Mathematicians will have you do weird definitions and waste their time doing this pointless formalism where perhaps I doubt that this has an inverse when it seems as though it obviously does. lol

OpenStudy (anonymous):

i thought u only use the horizontal line test..

ganeshie8 (ganeshie8):

you're right, its horizontal test... lets go thru Kai

OpenStudy (anonymous):

ohk so what ur saying is that you can try to find the inverse, and than use f(f^−1(x))=x to check id its true?

OpenStudy (anonymous):

i mean f(f^-1(x))=x

ganeshie8 (ganeshie8):

yes if its easy, however perl's statement works in general : `f'(x) doesn't change sign ==> f(x) has an inverse`

OpenStudy (kainui):

Yeah that was more opinion and me being critical about the current style that mathematics has since I don't find the concept of injective, surjective, bijective, sets, etc useful or appealing or meaningful to me at all. If at any point your function is multivalued you can always just define an inverse for each part depending on which is meaningful to you. I can't see what you wrote @swift_13 it looks like a bunch of black question marks to me.

OpenStudy (anonymous):

if f(x)=1/x and its inverse is also f(x)^-1=1/x

ganeshie8 (ganeshie8):

f(x) = 1/x f'(x) = -1/x^2 which is always negative no matter what the x value is, so the graph is always decreasing which clearly tells you that it passes the horizontal line test

OpenStudy (anonymous):

and f(f(x)^-1)=x

ganeshie8 (ganeshie8):

if f(x) is a complicated function, you don't need to find the inverse to know whether an inverse exists or not

OpenStudy (anonymous):

oh, ohk cause i was just going through my text book and it just mentions the f(x)' >0 to have an inverse

OpenStudy (anonymous):

nothing about f(x)'<0

ganeshie8 (ganeshie8):

maybe take a screenshot and attach..

ganeshie8 (ganeshie8):

im not able to find a good sourse on this topic... but it makes sense intuituvely why the first derivative stuff works, right ?

OpenStudy (perl):

If f ' (x) < 0 on the interval, then f is decreasing on the interval. If f is decreasing on the interval , then f is monotone. Then use this theorem I paste here. https://proofwiki.org/wiki/Inverse_of_Strictly_Monotone_Function (read the part about inverse)

OpenStudy (anonymous):

yes it does make sense, so just to make sure if either the function is increasing or decreasing than it has an inverse

OpenStudy (jhannybean):

First time I'm hearing of a function being monotone, haha.

OpenStudy (perl):

lol

OpenStudy (anonymous):

so if f(x) = x^2 -2x

OpenStudy (anonymous):

f(x)'=2x

OpenStudy (jhannybean):

f'(x) = 2x - 2 *

OpenStudy (perl):

monotone functions have two possible states, they are either decreasing or they are increasing on the interval, but not both.

OpenStudy (perl):

that cant have an inverse because the derivative changes from negative to positive

OpenStudy (anonymous):

sorry yyea f(x)'=2x -2 thats means the function is both increasing and decreasing right

OpenStudy (perl):

correct

OpenStudy (anonymous):

so it can't have an inverse

OpenStudy (anonymous):

thanks a lot!

OpenStudy (perl):

no problem

OpenStudy (anonymous):

also i kind of forgot but how do u check if the function is strictly increasing or decreasing or both

OpenStudy (jhannybean):

And the function itself does not pass the horizontal line test.

OpenStudy (anonymous):

i mean its simple to c from f(x)'=2x -1 but for other functions

OpenStudy (anonymous):

the only way to check for horizontal line test is to draw a graph right? another way?

ganeshie8 (ganeshie8):

you have derivative now

ganeshie8 (ganeshie8):

f(x) passes horizontal line test if : f'(x) > 0 for all x or f'(x) < 0 for all x

ganeshie8 (ganeshie8):

its an exclusive or, not just or

OpenStudy (jhannybean):

Yeah, and because your derivative \(\ f'(x) = 2x - 2\) moves from negative values to positive values, it fails the test.

ganeshie8 (ganeshie8):

you could also try proving \(\large f(b) \gt f(a)\) whenever \(\large b \gt a\) or vice versa

ganeshie8 (ganeshie8):

above trick is more used in sequences though..

OpenStudy (anonymous):

ohk thanks guys!

OpenStudy (anonymous):

but like for f(x)'=2x-2 , would i say for f(x)' > 0 for all x>1 and f(x)'<0 for all x <1

OpenStudy (anonymous):

due to this the function doesn't have an inverse

OpenStudy (anonymous):

i mean due to this it doesn't pass the horizontal line test

OpenStudy (perl):

f '(x) <0 for x<1

OpenStudy (perl):

thats correct, the derivative has to stay positive or stay negative in order for the inverse of the original function to exist

ganeshie8 (ganeshie8):

yes.. however if you take only the piece x>1, then it surely has an inverse because that piece alone passes the horizontal line test.. but it would be like chaning the original problem... so you simply say : `the given function has no inverse because f'(x) changes signs`

OpenStudy (anonymous):

yup thanks!

OpenStudy (anonymous):

yup got it! i understand now!

OpenStudy (jhannybean):